# Large Language Model

### v1

@@ 12 Discussion
Training Methods. In this section, we discuss the various training approaches employed in our methods, including full fine-tuning, LoRA [19], and LLM Adaptors [20]. LoRA is commonly used in contexts with limited computational resources and small datasets. We initially experimented with LoRA for training our framework on a small batch of data, but found that its multiple manually-tuned parameters increased the risk of training instability within our framework. We also opted not to use adaptor-based training, as adaptors consist of learnable parameters tied to the original modality, which can limit model performance. Event data, containing more comprehensive information about the captured scenes than traditional camera data [37], requires a large parameter space to adequately capture this information. Given our access to abundant, high-quality paired data, we selected full fine-tuning as the primary training method for our framework. This approach effectively captures event-based categorical knowledge, supporting a deeper understanding of event-based world semantics.
@@
Scale. Our processed datasets span a variety of domains, including static images [23, 31], body motion [40], and driving scenes [16]. However, the data remains insufficient in certain areas, such as sports movements and complex scenes, which limits the model’s depth of understanding of event data. This is particularly evident in the driving domain, where high-resolution event data is still lacking, hindering the clarity of perception and collaborative development between event and regular cameras. To address this, we aim to construct a larger and more comprehensive dataset. Additionally, due to computational resource limitations, we are constrained to using models with fewer than 3 billion parameters for training or fine-tuning. We believe that our proposed framework can be readily scaled to models with a larger parameter count, enhancing its capacity for event-based understanding.
@@
Event Representation. In line with the current VLM setting [46, 48, 25], we use a red-blue color map event representation for processing and training. While this approach is effective, it may not be optimal for an MLLM-based framework. In future work, we plan to further investigate the impact of different event representations [41, 50] to explore alternatives that may better support our framework’s performance.

### v2

@@ Appendix E Discussion
Training Methods. In this section, we discuss the various training approaches employed in our methods, including full fine-tuning, LoRA [53], and LLM Adaptors [49]. LoRA is often used in scenarios with limited computational resources and small datasets. Initially, we experimented with LoRA for training our framework on a small batch of data, but found that its multiple manually-tuned parameters increased the risk of training instability. Additionally, we opted not to use adaptor-based training, as adaptors consist of learnable parameters tied to the original modality, which can constrain model performance. Event data, which contains more comprehensive information about the captured scenes compared to traditional camera data [23], requires a larger parameter space to fully capture this information. Given our access to abundant, high-quality paired data, we selected full fine-tuning as the primary training method for our framework. This approach effectively captures event-based categorical knowledge and supports a deeper understanding of event-based world semantics.
@@
Scale. Our processed datasets span a variety of domains, including static images [15, 26], body motion [16], and driving scenes [25]. However, certain areas, such as sports movements and complex scenes, remain underrepresented, limiting the model’s depth of understanding of event data. This is particularly evident in the driving domain, where high-resolution event data is still lacking, hindering both the clarity of perception and the collaborative development between event and regular cameras. To address this, we aim to construct a larger and more comprehensive dataset. Additionally, due to computational resource constraints, we are currently limited to using models with fewer than 3 billion parameters for training or fine-tuning. However, we believe that our proposed framework can easily scale to models with larger parameter counts, thereby enhancing its capacity for event-based understanding.
@@
Event Representation. In line with current VLM settings [22, 8, 9], we use a red-blue color map event representation for processing and training. While this approach has proven effective, it may not be optimal for an MLLM-based framework. In future work, we plan to investigate the impact of different event representations [54, 55] and explore alternatives that could better support our framework’s performance.